OpenStudy (anonymous):

convert 33.4% to a fraction in the lowest terms

OpenStudy (mertsj):

\[33.4 percent = .334=\frac{334}{1000}=\frac{167}{500}\]

OpenStudy (rulnick):

33.4/100 = 334/1000 = 167/500.

OpenStudy (anonymous):

Thank you

OpenStudy (mertsj):

yw